Q3. Suppose that

C = 60 + 0.8YD

I = 150 - 10r

G = 250

T = 200

Ms = 100

Md = 40 + 0.1Y - 10r

( all the magnitudes are in real terms)

a. Write the equations for the IS and LM schedules.

b. Find the equilibrium values for income and the interest rate .

Suppose we change the model such that investment is assumed to be completely

interest inelastic; investment does not depend on the rate of interest and we have

I = 150.

c. Write the new equations for the IS and LM schedules. Show the schedules graphically.

d. Find the new equilibrium values for income and the interest rate. (4+2+8+3=17 marks)


1
Expert's answer
2021-06-22T13:33:07-0400

a) "IS:\ I=Y-C-G+T,\ 150-10r=Y-60-0.8Y-250+200"

"Y=1300-50r"

LM:

"Ms=Md"

"100=40+0.1Y-10r"

"Y=600+100r"

b) "1300-50r=600+100r"

"r=4\frac{2}{3}=466.67\%" - equilibrium; "Y=1066.67" - equilibrium

c and d) IS:

"150=Y-60-0.8Y-250+200"

"Y=1300" - equilibrium

LM:

"100=40+0.1Y-10r"

"Y=600+100r"

"1300=600+100r"

"r=7=700\%" - equilibrium